# Unitary coupled-cluster singles and doubles (UCCSD) calculations in conjunction with fragment molecular orbital (FMO) scheme Kenji Sugisaki , Tatsuya Nakano, Yuji Mochizuki

### 28 Feb 2024 v1

@@ Abstract
The fragment molecular orbital (FMO) method is one of the popular methods to efficiently treat macromolecular systems by dividing the system of interest into small fragments based on embedding under the electrostatic potential (ESP). Such a fragmentation method has the potential advantage of making the circuit flat in quantum chemical calculations on quantum computers. In this study, we used a GPU-accelerated quantum simulator (cuQuantum) to perform the electron correlation part of the FMO calculation as a unitary coupled-cluster singles and doubles (UCCSD) with the variational quantum eigensolver (VQE), using hydrogen-bonded systems (FH) and (FH)-HO as testbeds. From the numerical simulations, we found that the Trotter decomposition of the UCCSD ansatz breaks the orbital-invariance condition, and it can yield different total energies for symmetrically equivalent molecules. We also observed that the Trotterized UCCSD does not satisfy the size-consistency, which is an essential requirement in the FMO scheme, when the molecular orbitals delocalized to dimers are used. The GPU acceleration was substantial for the simulations with larger numbers of qubits, and it was about a factor of 5.8–7.5 for 18 qubit systems.
@@
Fragment Molecular Orbital (FMO), Unitary Coupled-Cluster (UCC), cuQuantum, GPU, Size-Consistency, Orbital-Invariance, Trotter Error
@@ 1. Introduction
Starting with the seminal work of Aspuru-Guzik et al. [1], quantum chemical computation has been actively explored and developed as a promising application area for quantum computers [2, 3, 4, 5, 6], where the potential applicability to huge-scale configuration interactions [7] is attractive. In practice, however, the development of computational methods and algorithms using quantum simulators is currently more mainstream than the use of actual devices. For noisy intermediate-scale quantum (NISQ) computers, the unitary coupled-cluster singles and doubles (UCCSD) [8, 9] has been used for relatively small molecules in conjunction with the variational quantum eigensolver (VQE) [10, 11, 12, 13], and this VQE-UCCSD scheme has been extended to multi-reference cases, e.g., Refs. [14, 15, 16, 17]. In addition, GPU-accelerated simulators such as cuQuantum [18] have attracted considerable interest due to its pronounced performance [19].
In another direction, the so-called problem decomposition approach has been introduced to shallow the circuit depth [20] while avoiding the effects of noise. Note that such an approach is rather common for large molecules (like proteins) as the fragmentation-based methods [21, 22, 23]. The introduction of problem decomposition to quantum computation was pioneered by Yamazaki et al [24], who compared three methods of fragment molecular orbital (FMO) [25], divide-and-conquer (DC) [26], and density matrix embedding theory (DMET) [27]. Recently, the FMO calculations with VQE-UCCSD for hydrogen clusters have been reported [28].
In this study, we applied the VQE-UCCSD scheme [10, 11, 12, 13] to compute the electron correlation part of FMO calculations [25, 29, 30] for a couple of hydrogen-bonded systems, (FH) and (FH)-HO. For execution, the cuQuantum simulator [18] was used as done in the previous study [19]. In the following we describe the calculation method, the results and the discussion.
@@ 2.1. FMO scheme and program
Electron correlation calculations, such as second-order Møller–Plesset perturbations (MP2) [31], are performed after the HF calculations for each monomer are complete and after the individual HF calculations for each dimer are complete. The correlation energy correction is done in an additive manner as in Eq. (1). The introduction of electron correlations is essential to improve quantitatively by incorporating dispersion stabilization and reducing excess ionicities of the HF description. As described in Ref. [31], both size-consistency and orbital-invariance are crucial requirements in the correlated methods. This is obviously true for the FMO scheme based on Eq. (1).
Currently, GAMESS-US [32], PAICS [33], and ABINIT-MP [34, 35] are the available programs that can perform FMO calculations including electron correlation correction by MP2. Besides the MP2 capability [36, 37, 38], ABINIT-MP is unique in supporting higher correlated calculations [31, 39] on-the-fly; from the third-order MP (MP3) [40] to coupled-cluster singles and doubles including perturbative triples (CCSD(T)) [41] are supported.
@@ 2.2. Preparation of molecular integrals under FMO scheme
The geometries of (FH) (under C symmetry) and (FH)-HO (C symmetry) were optimized by the GAUSSIAN16W program [42] at the level of B3LYP [43] corrected with the empirical dispersion [44] with the 6-31+G(d’,p’) basis set [45]. The resulting Cartesian coordinates are listed in Table 1 and illustrated in Figure 1.
@@
For (FH) and (FH)-HO, the FMO calculations were performed with the STO-3G minimal basis set [46], where we used a local version of ABINIT-MP, which dumped the integral list of basis functions and the converged MO coefficients (at the FMO-HF level) of monomers and dimers as separate files. These data were transformed by a small Fortran program into molecular integrals for the second-quantized Hamiltonian used to run VQE-UCCSD, expressed as
Indices of and cover the correlating orbital space, and and are the transformed one- and two-electron integrals. Note that there is some degree of locality of monomer orbitals in dimer orbitals with respect to the occupied space for (FH) and also that there is the symmetric delocalization for the FH dimer in (FH)-HO. Due to a proof-of-concept (PoC) phase of this study, the FMO calculations by ABINIT-MP were done in a separate step from the quantum calculations described in the next subsection. The 1s orbitals of fluorine and oxygen were frozen [47] for in Eq. (2), which is a good approximation to save on the number of qubits [48, 49]. The number of correlated electrons for dimers was thus 16. For comparison with the VQE-UCCSD correlation energies, the usual FMO-MP2 and FMO-CCSD(T) calculations were also performed by ABINIT-MP. These calculations were completed in less than 1 second on a single core of Intel Xeon processor.
@@ 2.3. Set-up of VQE-UCCSD calculation
Various types of ansatzes have been proposed and studied for quantum chemical calculations [50]. Here, we adopted the UCCSD ansatz defined in Eqs. (3) and (4), because it is a chemically motivated ansatz and it can give very accurate correlation energies.
@@
In the implementation of the UCCSD quantum circuit, we adopted the first-order Trotter decomposition with one Trotter slice in conjunction with the magnitude ordering [54] of the cluster operators. For the variational optimization of the excitation amplitudes, we examined COBYLA [55] and Powell [56] algorithms; the corresponding labels are shortly denoted as UCCSD:CB and UCCSD:PW, respectively (see Tables 2 and 3). For comparison, the calculation of the complete active space configuration interaction (CAS-CI) was performed to obtain the exact correlation energy in the orbital space of STO-3G. The numerical simulations for (FH) and (FH)-HO were carried out on the Supercomputer ‘Flow’ Type-II subsystem at Nagoya University and on the in-house NVIDIA DGX H100 system, respectively.
@@ 3.1. Energies and timings
The correlation energies for (FH) are summarized in Table 2. Compared to MP2, CCSD has a significantly lower energy, and CCSD(T) gives values close to CAS-CI, as expected. The quantum calculation for UCCSD is obtained by two different solvers, COBYLA and Powell. Powell gave lower energies than COBYLA, but the number of function evaluations (total energy calculations) required in Powell is about 1.6–2.8 times greater than in COBYLA (115, 130, 119, 1762, 1963, and 1910 for “1”, “2”, “3”, “21”, “31”, and “32”, respectively, for COBYLA, and 225, 257, 193, 4783, 4765, and 5408 for Powell).
| Unit | MP2 | CCSD | CCSD(T) | UCCSD:CB | UCCSD:PW | CAS-CI |
@@
| “21” | 0.035493 | 0.051856 | 0.051933 | 0.049880 | 0.050452 | 0.051963 |
| “31” | 0.035980 | 0.052778 | 0.052850 | 0.051554 | 0.051527 | 0.052879 |
| “32” | 0.035446 | 0.053122 | 0.053123 | 0.051952 | 0.053067 | 0.053124 |
| Sum. | 0.053527 | 0.077666 | 0.077816 | 0.073439 | 0.075065 | 0.077876 |
Table 3 summarizes the results for the (FH)-HO correlation energies. A checkpoint here is whether the equivalence symmetries (monomers “2” and “3” / dimers “21” and “31”) are satisfied, and the usual MP2, CCSD, CCSD(T), and CAS-CI results all satisfy this requirement. The trend of the correlation energies by these methods is the same as in (FH). On the other hand, the UCCSD results (of both COBYLA and Powell) unfortunately do not satisfy symmetry, as the difference is seen to five decimal places for monomers and three decimal places for dimers. From a chemical precision point of view, it seems problematic that the effect is seen to three decimal places. Furthermore, this issue of broken equivalence should be kept in mind not only for FMO, but for all approaches of fragmentation-oriented methods [21, 22, 23]. This problem of VQE-UCCSD is related to the Trotter error and is discussed in the next section.
| Unit | MP2 | CCSD | CCSD(T) | UCCSD:CB | UCCSD:PW | CAS-CI |
| Monomer | ||||||
| “1” | 0.035370 | 0.049321 | 0.049394 | 0.049242 | 0.049231 | 0.049445 |
| “2” | 0.017810 | 0.026705 | 0.026705 | 0.026608 | 0.026692 | 0.026705 |
| “3” | 0.017810 | 0.026705 | 0.026705 | 0.026677 | 0.026659 | 0.026705 |
| Dimer | ||||||
| “21” | 0.053486 | 0.075392 | 0.075526 | 0.069313 | 0.071421 | 0.075600 |
| “31” | 0.053486 | 0.075392 | 0.075526 | 0.071847 | 0.072544 | 0.075600 |
| “32” | 0.035790 | 0.053549 | 0.053553 | 0.053230 | 0.053207 | 0.053549 |
| Sum. | 0.071770 | 0.101602 | 0.101801 | 0.091863 | 0.094590 | 0.101895 |
The cuQuantum quantum simulator was used in this VQE-UCCSD computation. Table 4 summarizes the timings of the UCCSD jobs of (FH) on the ‘Flow’ Type-II subsystem with and without GPU. The GPU acceleration was about a factor of 1.8–2.5 and 5.8–7.5 for monomers and dimers, respectively. GPU acceleration is substantial for larger systems, but the speedup is less significant compared with our previous study [19]. This is because the VQE job needs a lot of time for pre-processing and post-processing, and these parts cannot be accelerated by cuQuantum. Considering that a normal FMO-CCSD(T)/STO-3G calculation takes less than 1 second to complete, there is a speed difference of the order of the fourth power of 10 if the correlation part is due to VQE-UCCSD at this time. Note that the present VQE-UCCSD was run on a classical computer, where computational time grows exponentially with the number of qubits. Anyway, as in the previous report [19], GPU acceleration with cuQuantum is essential for quantum simulations.
| UCCSD:CB | UCCSD:PW | |||||
| Unit | with GPU | without GPU | Acceleration | with GPU | without GPU | Acceleration |
| Monomer | ||||||
| “1” | 14.1 | 30.1 | 2.13 | 25.4 | 52.3 | 2.06 |
| “2” | 16.1 | 29.1 | 1.80 | 27.9 | 68.9 | 2.47 |
| “3” | 14.8 | 31.5 | 2.13 | 22.2 | 48.4 | 2.18 |
| Dimer | ||||||
| “21” | 11901.3 | 70531.8 | 5.93 | 31276.4 | 216982.7 | 6.94 |
| “31” | 12871.9 | 74618.2 | 5.80 | 30315.4 | 210808.9 | 6.95 |
| “32” | 11209.1 | 83915.6 | 7.49 | 27740.4 | 203298.3 | 7.33 |

### 13 Mar 2024 v2

# Size-consistency and orbital-invariance issues revealed by VQE-UCCSD calculations with the FMO scheme Kenji Sugisaki , Tatsuya Nakano, Yuji Mochizuki

@@ Abstract
The fragment molecular orbital (FMO) scheme is one of the popular fragmentation-based methods and has the potential advantage of making the circuit flat in quantum chemical calculations on quantum computers. In this study, we used a GPU-accelerated quantum simulator (cuQuantum) to perform the electron correlation part of the FMO calculation as unitary coupled-cluster singles and doubles (UCCSD) with the variational quantum eigensolver (VQE) for hydrogen-bonded (FH) and (FH)-HO systems with the STO-3G basis set. VQE-UCCD calculations were performed using both canonical and localized MO sets, and the results were examined from the point of view of size-consistency and orbital-invariance affected by the Trotter error. It was found that the use of localized MO leads to better results, especially for (FH)-HO. The GPU acceleration was substantial for the simulations with larger numbers of qubits, and was about a factor of 6.7–7.7 for 18 qubit systems.
@@
FMO, UCC, GPU, VQE, Trotter Error
@@ 1. Introduction
Starting with the seminal work of Aspuru-Guzik et al. [1], quantum chemical computation has been actively explored and developed as a promising application area for quantum computers [2, 3, 4, 5, 6], where the potential applicability to huge-scale configuration interactions such as the FeMo-cofactor of nitrogenase [7] is attractive with care for the setting of active orbital space [8]. In practice, however, the development of computational methods and algorithms using quantum simulators is currently more mainstream than the use of actual devices. For noisy intermediate-scale quantum (NISQ) computers, the unitary coupled-cluster singles and doubles (UCCSD) [9, 10, 11, 12, 13, 14, 15, 16] has been used for relatively small molecules in conjunction with the variational quantum eigensolver (VQE) [17, 18, 19, 20], and this VQE-UCCSD scheme has been extended to multi-reference cases, e.g., Refs. [21, 22, 23, 24]. In addition, GPU-accelerated simulators e.g., cuQuantum [25] have attracted considerable interest due to its pronounced performance [26].
In another direction, the so-called problem decomposition approach has been introduced to shallow the circuit depth [27] while avoiding the effects of noise. Note that such an approach is rather common for large molecules (like proteins) as the fragmentation-based methods [28, 29, 30]. The introduction of problem decomposition to quantum computation was pioneered by Yamazaki et al. [31], who compared three methods of fragment molecular orbital (FMO) [32], divide-and-conquer (DC) [33], and density matrix embedding theory (DMET) [34]. Recently, the FMO calculations with VQE-UCCSD for hydrogen clusters have been reported [35]. Note that the (H), trans-butadiene, and [Cr(OH)(NH)] systems were treated by UCCSD based on the concept of orbital locality [36].
In this study, we applied the VQE-UCCSD scheme [17, 18, 19, 20] to compute the electron correlation part of FMO calculations [32, 37, 38] for a couple of hydrogen-bonded systems, (FH) and (FH)-HO. For execution, the cuQuantum simulator [25] was used as done in the previous study [26]. Effect of Trotterization on the orbital-invariance condition of the UCCSD method was investigated using two symmetrically equivalent FH molecules in the latter system. We also studied relationship between the size-consistency condition and the Trotterized UCCSD ansatz, using square tetrahydrogen (4H) and cuboid octahydrogen (8H) clusters. Acceleration of the VQE-UCCSD simulations using cuQuantum is also discussed. The rest of the paper is organized as follows: In Section 2, we describe the calculation methods of both FMO stage and VQE-UCCSD stages. The results of the FMO correlation energies are shown first, and then the issues surrounding the Trotter error are discussed in Section 3. In Section 4, we summarize our work and discuss possible directions for future work.
@@ 2.1. FMO scheme and program
Electron correlation calculations, such as second-order Møller–Plesset perturbation (MP2) [39], are performed after the HF calculations for each monomer are complete and after the individual HF calculations for each dimer are complete. The correlation energy correction is done in an additive manner as in Eq. (1). The introduction of electron correlations is essential to improve quantitatively by incorporating dispersion stabilization and reducing excess ionicities of the HF description. As described in Ref. [39], both size-consistency and orbital-invariance are crucial requirements in the correlated methods. This is obviously true for the FMO scheme based on Eq. (1).
Currently, GAMESS-US [40, 41], PAICS [42, 43], and ABINIT-MP [44, 45] are the available programs that can perform FMO calculations including electron correlation correction by MP2. Besides the MP2 capability [46, 47, 48], ABINIT-MP is unique in supporting higher-order correlated calculations [39, 49] on-the-fly; from the third-order MP (MP3) [50] to coupled-cluster singles and doubles including perturbative triples (CCSD(T)) [51] are supported.
@@ 2.2. Preparation of molecular integrals under FMO scheme
The geometries of (FH) (under C symmetry) and (FH)-HO (C symmetry) were optimized by the GAUSSIAN16W program [52] at the level of B3LYP [53] corrected with the empirical dispersion [54] with the 6-31+G(d’,p’) basis set [55]. The resulting Cartesian coordinates are listed in TABLE 1 and illustrated in FIG. 1.
@@
For (FH) and (FH)-HO, the FMO calculations were performed with the STO-3G minimal basis set [56], where we used a local version of ABINIT-MP, which dumped the integral list of basis functions and the converged canonical MO (CMO) coefficients (at the FMO-HF level) of monomers and dimers as separate files. These data were transformed by a small Fortran program into molecular integrals for the second-quantized Hamiltonian used to run VQE-UCCSD, expressed as
Indices of and cover the correlating orbital space, and and are the transformed one- and two-electron integrals. The 1s-like CMOs of fluorine and oxygen were frozen [57] for in Eq. (2), which is a good approximation to save on the number of qubits [58, 59]. The number of correlated electrons for dimers was thus 16.
Note that there is some degree of locality of monomer CMOs in dimer orbitals with respect to the occupied space for (FH) and also that there is the symmetric delocalization for the FH dimer in (FH)-HO. To address the issue of size-consistency, the Pipek–Mezey localization[60] was performed for the valence occupied CMOs and the virtual CMOs, respectively, and these sets of localized MOs (LMOs) were also used for the integral transformation. The lists of molecular orbitals (CMOs and LMOs) of for the monomers and dimers of (FH) and (FH)-HO are shown in FIG. S1 and S2, respectively, in Supplementary Materials.
Due to a proof-of-concept (PoC) phase of this study, the FMO calculations (at the HF level) by ABINIT-MP were done in a separate step from the quantum calculations described in the next subsection. For comparison with the VQE-UCCSD correlation energies, the usual FMO-MP2 and FMO-CCSD(T) calculations were also performed by ABINIT-MP. These calculations were completed in less than 1 second on a single core of Intel Xeon processor.
@@ 2.3. Set-up of VQE-UCCSD calculation
Various types of ansatzes have been proposed and studied for quantum chemical calculations [61]. In this work, we adopted the UCCSD ansatz defined in Eqs. (3) and (4), because it is a chemically motivated ansatz and it can give very accurate correlation energies.
@@
In the implementation of the UCCSD quantum circuit, we adopted the first-order Trotter decomposition given in Eq. (5) in conjunction with the magnitude ordering [65] of the cluster operators.
Here, is the excitation/de-excitation operators in the Pauli operator expressions obtained by adopting the SCBKT to the operator in the second-quantized form. is a direct product of Pauli operators called as a Pauli string, and is the corresponding coefficient derived from and . is the number of Pauli strings, and is the number of Trotter slices. Unless otherwise specified we used the one Trotter slice () for the VQE-UCCSD simulations.
For the variational optimization of the excitation amplitudes, we examined COBYLA [66] and Powell [67] algorithms; the corresponding labels are shortly denoted as UCCSD:CB and UCCSD:PW, respectively (see TABLE 2 and 3). For comparison, the calculation of the complete active space configuration interaction (CAS-CI) was performed to obtain the exact correlation energy in the orbital space of STO-3G. The numerical simulations for (FH) and (FH)-HO were carried out on the Supercomputer ‘Flow’ Type-II subsystem at Nagoya University and on the in-house NVIDIA DGX H100 system, respectively.
@@ 3.1. Energies and timings
The correlation energies for (FH) are summarized in TABLE 2. Compared to MP2, CCSD has a significantly lower energy, and CCSD(T) gives values close to CAS-CI, as expected. UCCSD:PW gave lower energies than UCCSD:CB, but the number of function evaluations (total energy calculations) required in Powell is about 1.6–2.8 times greater than in COBYLA (see TABLE S1 in Supplementary Materials for details). The same trend was observed for the LMO-based UCCSD calculations. No significant difference was found in the number of function evaluations between CMO and LMO-based calculations.
As we discuss in the next section, Trotterized UCCSD does not automatically satisfy the size-consistency condition, and using LMOs as the basis is crucial to ensure that Trottterized UCCSD is size-consistent. In fact, the correlation energies of the dimers are improved in the LMO-based UCCSD:PW calculations, and the sum of the correlation energies is 0.001679 Hartree (1.0536 kcal mol) lower in the LMO-based calculations than in the CMO-based one. The deviation of the sum of UCCSD:PW correlation energy from the CAS-CI one is 0.71 kcal mol.
| Unit | MP2 | CCSD | CCSD(T) | UCCSD:CB | UCCSD:PW | CAS-CI | ||
| CMO333HF canonical orbitals were used for the calculation. | LMO444Localied molecular orbitals constructed by using Pipek–Mezey method were used for the calculation. | CMO | LMO | |||||
| Monomer | ||||||||
| “1” | 0.017933 | 0.026945 | 0.026945 | 0.026884 | 0.026854 | 0.026914 | 0.026729 | 0.026945 |
| “2” | 0.017526 | 0.026216 | 0.026216 | 0.026164 | 0.026169 | 0.026192 | 0.026019 | 0.026216 |
| “3” | 0.017933 | 0.026929 | 0.026929 | 0.026899 | 0.026839 | 0.026875 | 0.026691 | 0.026929 |
| Dimer | ||||||||
| “21” | 0.035493 | 0.051856 | 0.051933 | 0.049880 | 0.050429 | 0.050452 | 0.051322 | 0.051963 |
| “31” | 0.035980 | 0.052778 | 0.052850 | 0.051554 | 0.051632 | 0.051527 | 0.052169 | 0.052879 |
| “32” | 0.035446 | 0.053122 | 0.053123 | 0.051952 | 0.051317 | 0.053067 | 0.052692 | 0.053124 |
| Sum. | 0.053527 | 0.077666 | 0.077816 | 0.073439 | 0.073516 | 0.075065 | 0.076744 | 0.077876 |
TABLE 3 summarizes the results for the (FH)-HO correlation energies. The number of function evaluations in the VQE-UCCSD optimization is given in TABLE S2 in Supplementary Materials. A checkpoint here is whether the equivalence symmetries (monomers “2” and “3” / dimers “21” and “31”) are satisfied, and the usual MP2, CCSD, CCSD(T), and CAS-CI results all satisfy this requirement. The trend of the correlation energies by these methods is the same as in (FH). On the other hand, the UCCSD results (of both COBYLA and Powell) unfortunately do not satisfy symmetry, as the difference is seen to five decimal places for monomers and three decimal places (in the order of kcal mol) for dimers. From a chemical precision point of view, it seems problematic that the effect is seen to three decimal places. Furthermore, this issue of broken equivalence should be kept in mind not only for FMO, but for all approaches of fragmentation-oriented methods [28, 29, 30]. This problem of VQE-UCCSD is related to the Trotter error and is discussed in the next section.
The effect of orbital localization on the UCCSD correlation energy is remarkable in the (FH)-HO system. In the UCCSD:PW calculations the sum of correlation energies improved about 0.005256 Hartree (3.2982 kcal mol) by the orbital localization, and deviation from the CAS-CI correlation energy is 1.29 kcal mol. Note that orbital localization also improves the orbital-invariance condition. By using the LMOs, the difference in correlation energy between “21” and “31” is reduced to 0.039 kcal mol. These results exemplify the importance of using LMOs in the combination of FMO and VQE-UCCSD approaches.
| Unit | MP2 | CCSD | CCSD(T) | UCCSD:CB | UCCSD:PW | CAS-CI | ||
| CMO777HF canonical orbitals were used for the calculation. | LMO888Localied molecular orbitals constructed by using Pipek–Mezey method were used for the calculation. | CMO | LMO | |||||
| Monomer | ||||||||
| “1” | 0.035370 | 0.049321 | 0.049394 | 0.049242 | 0.049214 | 0.049231 | 0.048915 | 0.049445 |
| “2” | 0.017810 | 0.026705 | 0.026705 | 0.026608 | 0.026648 | 0.026692 | 0.026538 | 0.026705 |
| “3” | 0.017810 | 0.026705 | 0.026705 | 0.026677 | 0.026637 | 0.026659 | 0.026476 | 0.026705 |
| Dimer | ||||||||
| “21” | 0.053486 | 0.075392 | 0.075526 | 0.069313 | 0.071305 | 0.071421 | 0.074429 | 0.075600 |
| “31” | 0.053486 | 0.075392 | 0.075526 | 0.071847 | 0.069583 | 0.072544 | 0.074367 | 0.075600 |
| “32” | 0.035790 | 0.053549 | 0.053553 | 0.053230 | 0.052161 | 0.053207 | 0.052979 | 0.053549 |
| Sum. | 0.071770 | 0.101602 | 0.101801 | 0.091863 | 0.090551 | 0.094590 | 0.099846 | 0.101895 |
The cuQuantum quantum simulator was used in this VQE-UCCSD computation. TABLE 4 summarizes the timings of the UCCSD jobs of (FH) using LMOs on the ‘Flow’ Type-II subsystem with and without GPU. The GPU acceleration was about a factor of 1.6–2.2 and 6.7–7.7 for monomers and dimers, respectively. The VQE-UCCSD simulations of the dimers “21” and “31” of (FH)-HO (20 qubit systems) without GPU acceleration are too time-consuming to do. Here we estimated the acceleration ratio by performing the UCCSD simulations of the dimer “21” of (FH)-HO by setting the maximum number of function evaluations in the VQE parameter optimization to be 100 on NVIDIA DGX H100. With the GPU acceleration, the time required for pre-processing (Fermion–qubit transformation, reference CAS-CI calculation, MP2 and MP3 calculations, etc.) was 4773.9 seconds and 100 function evaluations in the VQE optimization took 1211.0 seconds. In contrast, the CPU-only calculation took 4421.2 and 37363.9 seconds for pre-processing and 100 function calls, respectively. The GPU acceleration of the VQE iteration part is about a factor of 30.85. Since the numbers of function evaluations required for convergence in UCCSD:CB and UCCSD:PW were 4913 and 12452, respectively, the time for the CPU-only simulations are estimated to be about 21 and 54 days for UCCSD:CB and UCCSD:PW, respectively. GPU acceleration is substantial for larger systems, but the speedup is less significant compared with our previous study [26]. This is because the VQE job needs a lot of time for pre-processing and post-processing, and these parts cannot be accelerated by cuQuantum. Considering that a normal FMO-CCSD(T)/STO-3G calculation takes less than 1 second to complete, there is a speed difference of the order of the fourth power of 10 if the correlation part is due to VQE-UCCSD at this time. Note that the present VQE-UCCSD was run on a classical computer, where computational time grows exponentially with the number of qubits. Anyway, as in the previous report [26], GPU acceleration with cuQuantum is essential for quantum simulations.
| Unit | UCCSD:CB | UCCSD:PW | ||||
| with GPU | without GPU | Acceleration | with GPU | without GPU | Acceleration | |
@@
| “21” | 12767.5 | 93205.0 | 7.30 | 34153.3 | 260996.7 | 7.64 |
| “31” | 13947.1 | 94199.7 | 6.75 | 34397.3 | 237695.3 | 6.91 |
| “32” | 8805.9 | 59975.6 | 6.81 | 21129.0 | 156257.2 | 7.40 |